Eksnr. 01 Kar. 5 Eksnr. 03 Kar. 5 Eksnr. 05 Kar. 10 Eksnr. 06 Kar. 11 Eksnr. 07 Kar. 5 Eksnr. 08 Kar. 00 Eksnr. 09 Kar. 03 Eksnr. 10 Kar. 00 Eksnr. 12 Kar. 6 Eksnr. 13 Kar. 10 Eksnr. 16 Kar. 11 Eksnr. 17 Kar. 11 Eksnr. 18 Kar. 7 Eksnr. 19 Kar. 5 Eksnr. 20 Kar. 11 Eksnr. 22 Kar. 10 Eksnr. 23 Kar. 00 Eksnr. 25 Kar. 8 Eksnr. 26 Kar. 03 Eksnr. 29 Kar. 7 Eksnr. 30 Kar. 00 Eksnr. 31 Kar. 7 Eksnr. 32 Kar. 10 Eksnr. 33 Kar. 8 Eksnr. 34 Kar. 7 Eksnr. 35 Kar. 7 Eksnr. 36 Kar. 03 Eksnr. 39 Kar. 11 Eksnr. 40 Kar. 10 Eksnr. 41 Kar. 9 Eksnr. 42 Kar. 7 Eksnr. 43 Kar. 03 Eksnr. 45 Kar. 10 Eksnr. 46 Kar. 03 Eksnr. 47 Kar. 9 Eksnr. 48 Kar. 5 Eksnr. 51 Kar. 8 Eksnr. 53 Kar. 10 Eksnr. 54 Kar. 7 Eksnr. 56 Kar. 6 Eksnr. 57 Kar. 5 Eksnr. 59 Kar. 5 Eksnr. 61 Kar. 8 Eksnr. 65 Kar. 10 Eksnr. 68 Kar. 03 Eksnr. 69 Kar. 9 Eksnr. 71 Kar. 10 Eksnr. 73 Kar. 9 Eksnr. 74 Kar. 00 Eksnr. 75 Kar. 7 Eksnr. 76 Kar. 5 Eksnr. 77 Kar. 13 Eksnr. 78 Kar. 10 Eksnr. 80 Kar. 7 Eksnr. 81 Kar. 10 Eksnr. 82 Kar. 5 Eksnr. 83 Kar. 7 Eksnr. 84 Kar. 5 Eksnr. 86 Kar. 7 Eksnr. 88 Kar. 00 Eksnr. 90 Kar. 6 Eksnr. 91 Kar. 9 Eksnr. 92 Kar. 00 Eksnr. 94 Kar. 8 Eksnr. 96 Kar. 8 Eksnr. 97 Kar. 8